OpenStudy (anonymous):

Is the system consistent or inconsistent? y = -1/2x - 1 and y = 3x + 3 Consistent Inconsistent Cannot be determined

OpenStudy (anonymous):

Surely A friend !:)

OpenStudy (anonymous):

Look : y=-1/2x-1 y=3x+3 ________________ -y=1/2x+1 y=3x+3 _____________ -y-1/2x=1 y-3x=3 ___________ -4/2x = 4 x= 4/-4/2 Now can you answer friend ?!:)

OpenStudy (anonymous):

consistent?

OpenStudy (anonymous):

@E.ali ???

OpenStudy (anonymous):

Yeah friend !

OpenStudy (anonymous):

thxx :))